The first term and the mth term of a geometrical progression are a and n respectively and its nth term is m .then its (m+1-n)th term is?

Grade:10

1 Answers

Shailendra Kumar Sharma
6 years ago
first term A=a
mth term
Ar(m-1) = n
nth term
Ar(n-1) = m
divide both 
r(m-n) =n/m
and (m-n+1)th term is Ar(m-n)  = a(n/m)